# On the adversarial robustness of Locality-Sensitive Hashing in Hamming space

### v2

@@ 5 Experiments
The goal of this section is to verify the efficacy of our adversary on real datasets, where an isolated point does not necessarily exist, and for parameter configurations outside of the guarantees of Theorems A.4 and A.6. The code can be found at https://github.com/makarov-mv/lsh-adversary.
Used resources.
The experiments were run on a server in an internal cluster, equipped with two AMD EPYC 7302 16-Core processors. All available cores were used. In total, running all of the experiments, including those in the appendix, takes several dozen hours, which is equivalent to hundreds of hours if they there run on a single core.
@@
Implementation details.
We use a custom implementation of LSH, which follows the approach of [${}^{+}$99]. It is the same as what was describe in Section 3 except that for each hash function’s buckets we don’t store all of the points from $P$ that hash into those buckets, but instead at most $B$ of them. In our implementation, we set $B=1$. More precisely, for each hash function we do a linear scan over all of the points. Each point is then put into its bucket only if no earlier point was already stored there. As for the implementation of the adversary, we use Algorithm 3 but consider the offset $t$ to be an input parameter instead of a fixed value. In each run in our experiments we reinitialize the LSH with new randomness, and if the used dataset is synthetic, we also generate a new one.
Experiment 1: Influence of parameters on success probability.
The goal of this experiment is to explore within which parameter setting the algorithm works. For this experiment we use synthetic datasets, since for them we can easily vary the number of points and dimensions. We use two datasets:
@@
Experiment 2: Efficiency across datasets.
In this experiment we measure how well our algorithm performs on a variety of different datasets. We use the datasets from Experiment 1, as well as one additional synthetic dataset and 3 datasets with real data.
The new synthetic dataset is Sparse random. Each point there is generated by independently sampling each bit to be $1$ with probability $1/15$, and to be 00 otherwise. The remaining datasets are:
MNIST [LeC98]. This dataset consists of grayscale pictures of handwritten digits of size $28\times 28$. We first flatten the images into vectors of size $784$, and then we project those vectors into Hamming space by setting each entry to 00 if it corresponded to a white pixel, i.e. if the grayscale color was 00, and to $1$ otherwise. Note that this transformation preserves the dataset fairly well, as the background pixels in each image are perfectly white.
Anonymous Microsoft Web Data (MSWeb) [BHK98]. Entries in this dataset are anonymous web user IDs and paths on the Microsoft website that they have visited. There are in total $294$ different paths. We represent each user by a point in $294$ dimensional Hamming space, where a point entry is equal to $1$ if they have visited the corresponding path.
Mushroom [UC 87, Wul82]. This dataset consists of descriptions of samples of mushrooms of $23$ different species. Each mushroom has $22$ categorical features describing it. We one-hot encode each feature to map the descriptions into Hamming space, with the final number of dimensions being $116$.
Those datasets are fixed in all of the experiments.
Due to computational constraints, we choose to limit the size of each dataset to at most $10000$. For MNIST and MSWeb we only use the first $10000$ points. We use all $8124$ points in Mushroom. For the synthetic datasets we generate $10000$ points with the dimension $300$, which is about the same order of magnitude as the dimensions of real datasets.
We choose a setting of parameters with which our algorithm has a good chance of success as per the results of Experiment 1. As the number of dimensions varies depending on the dataset, we set $r=0.15d$. $c=0$, $\lambda=4$ and the starting distance is 00 on the second subplot. Otherwise, we use the same setup as Experiment 1, except that for fixed datasets the origin point is the first one in the dataset.
On the first subplot the starting distance is measured as a fraction of $r$. On the second subplot the goal of the algorithm is to find a query to the LSH data structure that returns $\bot$ at a distance from the origin specified by the $x$ axis as a fraction of the $cr$. The dashed black vertical line represents the radius $r$ and is located at $0.5$ since $c=2$.
@@
The results in the first subplot of Figure 2 show that success probability drops monotonically for all datasets as the starting distance grows. This is natural to expect given that similar behaviour was observed in Experiment 1. On the second subplot we can see that for the most datasets we can reliably find a negative query only at distance much larger than $r$. This, in a way, allows us to compare the “hardness” of datasets, as further away from the origin we move, the less functions collide the origin point $z$ and our current point, the less bits we need to change for the query to become negative.
As one can see, all of the real datasets lie between the Random and Sparse random on the second subplot. This can be explained by the fact that the points in real datasets are much sparser than the Random, and the maximum distance from the most isolated point to the nearest point is smaller than the ideal $2cr$.
6 Conclusions
In this paper we have shown that it is in principle possible to attack standard LSH schemes and force false positives exponentially faster than it would be possible with non-adaptive queries. This underlines the importance of recent work on false-negative-free LSH [Pag18, Ahl17, Wei22] as well as adversarial robustness [${}^{+}$22, ACSS23, CNSS23] and motivates future research on LSH implementations that are both practically efficient and adversarial robust.
7 Discussion of impact
The main contribution of the paper is to get an improved understanding for which choices of parameters LSH data structure are unsafe wrt. adversarial attacks. We show that, in principle, attacks on standard LSH constructions are possible, if the parameters are not chosen carefully. As a consequence false-negative-free LSH data structures should be given preference in sensitive applications. Our main contribution is the theoretical analysis of the attack and our experiments are focussing on understanding the parameter setting for which attack may work. Therefore we believe that there is very limited risk to abuse our results and we do think that the benefit of understanding such potential attacks clearly outweights this limited risk.
@@ Appendix A Omitted proofs and statements
A.1 Proofs from Section 4.1
@@
Pr[|support of $\Pr\left[|\text{support of }g|<k-7\log n\cdot\max\left\{1,\frac{k(k-1)}{2d}\right\}\right]$
$=\Pr\left[S\geq k-7\log n\cdot\max\left\{1,\frac{k(k-1)}{2d}\right\}\right]$
$\leq\Pr\left[S^{\prime}\geq k-7\log n\cdot\max\left\{1,\frac{k(k-1)}{2d}\right\}\right]\leq 1/n^{3},$
@@ A.4 Existence of an isolated point
Appendix B Additional experiments
@@
In this experiment the goal was to measure the impact of $\lambda$ on the probability of success of the algorithm. The setup is the same as in Experiment 2, except $r=30$. All runs are done on the Zero dataset. The results are on Figure 3.
@@
Finally, we investigated the efficiency gain that our adaptive algorithm achieved over random point sampling. The experimental setup is the same as in Experiment 2, except $r=30$. Here we use the Random dataset. For both approaches, we continued rerunning them until a false negative query was found. The results are presented in Figure 5.

### v3

@@ 4.1 Analysis of the algorithm
5 Discussion of potential defenses against our attacking scheme
As was mentioned in Section 2, there already exists large literature on the robustification of LSH, many of which work against our adversary. In particular, we highlight [1], which provides a false-negative-free Las-Vegas type data structure with query time $dn^{1/c+o(1)}$ and used space $dn^{1+1/c+o(1)}$, achieving the same asymptotic performance as [22], as well as [14], which can be applied to LSH to guarantee perfect robustness by paying $O(d)$ times more space.
However, both approaches are not perfect within the same parameter bounds, as approach of [1] doesn’t give any guarantees on query time in this setting, and the space usage of [14] is too prohibitive in settings with large $d$. We experimentally investigate how the latter, and a construction based on differential privacy, fares against our adaptive adversary when the available space is limited in Experiment 6 in Appendix C. Interestingly, these defences perform well in this setting, which suggests that they might be reasonably effective even against a general adversary.
@@ 6 Experiments
The goal of this section is to verify the efficacy of our adversary on real datasets, where an isolated point does not necessarily exist, and for parameter configurations outside of the guarantees of Theorems A.4 and A.6.
@@
Implementation details. Our implementation of LSH follows Algorithm 1 in Section 3 [22]. To summarise, given a query point $q$, we iterate through all points in the dataset that hash together with $q$ under some hash function, then return the first one which lies at distance at most $cr$ from $q$, and return nothing if there is no such point. The adversary is implemented according to Algorithm 3.
Experiment 1: Influence of parameters on success probability. The goal of this experiment is to explore within which parameter setting the algorithm works. For this experiment we use synthetic datasets, since for them we can easily vary the number of points and dimensions. We use two datasets:
@@
Experiment 2: Efficiency across datasets. In this experiment we measure how well our algorithm performs on a variety of different datasets. We use the datasets from Experiment 1, as well as one additional synthetic dataset and 3 datasets with real data. The new synthetic dataset is Sparse random. Each point there is generated by independently sampling each bit to be $1$ with probability $1/15$, and to be 00 otherwise. The remaining datasets are: We use embeddings of the data sets MNIST [28], Anonymous Microsoft Web Data (MSWeb) [12] and Mushroom [34, 38] into Hamming space. For MNIST we flatten the images into vectors whose Boolean variables indicate the presence of a pixel. For MSWeb and Mushroom we use one-hot encoding to map the data into Hamming space. A more detailed description can be found in the Appendix C.
@@
Due to computational constraints, we choose to limit the size of each dataset to at most $10000$. For MNIST and MSWeb we only use the first $10000$ points. We use all $8124$ points in Mushroom. For the synthetic datasets we generate $10000$ points with the dimension $300$, which is about the same order of magnitude as the dimensions of real datasets. We choose a setting of parameters with which our algorithm has a good chance of success as per the results of Experiment 1. As the number of dimensions varies depending on the dataset, we set $r=0.15d$. $c=2$, $\lambda=4$ and the starting distance is 00 on the second subplot. Otherwise, we use the same setup as Experiment 1, except that for fixed datasets the origin point is the first one in the dataset. On the first subplot the starting distance is measured as a fraction of $r$. On the second subplot the goal of the algorithm is to find a query to the LSH data structure that returns $\bot$ at a distance from the origin specified by the $x$ axis as a fraction of the $cr$. The dashed black vertical line represents the radius $r$ and is located at $0.5$ since $c=2$. The results in the first subplot of Figure 2 show that success probability drops monotonically for all datasets as the starting distance grows. This is natural to expect given that similar behaviour was observed in Experiment 1. On the second subplot we can see that for the most datasets we can reliably find a negative query only at distance much larger than $r$. This, in a way, allows us to compare the “hardness” of datasets, as further away from the origin we move, the less functions collide the origin point $z$ and our current point, the less bits we need to change for the query to become negative. As one can see, all of the real datasets lie between the Random and Sparse random on the second subplot. This can be explained by the fact that the points in real datasets are much sparser than the Random, and the maximum distance from the most isolated point to the nearest point is smaller than the ideal $2cr$.
@@ Appendix A Omitted proofs and statements
A.1 Proofs from Section 4
@@
Pr[|support of $\Pr\left[|\text{support of }g|<k-7\log n\cdot\max\left\{1,\frac{k(k-1)}{2d}\right\}\right]\\ =\Pr\left[S\geq k-7\log n\cdot\max\left\{1,\frac{k(k-1)}{2d}\right\}\right]\\ \leq\Pr\left[S^{\prime}\geq k-7\log n\cdot\max\left\{1,\frac{k(k-1)}{2d}\right\}\right]\\ \leq 1/n^{3},$
@@ A.4 Existence of an isolated point
Appendix B Discussion of impact
The main contribution of the paper is to get an improved understanding for which choices of parameters LSH data structure are unsafe with respect to adversarial attacks. We show that, in principle, attacks on standard LSH constructions are possible, if the parameters are not chosen carefully. As a consequence false-negative-free LSH data structures should be given preference in sensitive applications. Our main contribution is the theoretical analysis of the attack and our experiments are focussing on understanding the parameter setting for which attack may work. Therefore we believe that there is very limited risk to abuse our results and we do think that the benefit of understanding such potential attacks clearly outweights this limited risk.
Appendix C Additional experiments and details
Resources used.
The experiments were run on a server in an internal cluster, where a typical worker has two AMD EPYC 7302 16-Core processors, with up to 500 GiB of RAM. Experiments where run in parallel on multiple cores. In total, running all of the experiments, including those in the appendix, takes at most 100000 CPU hours, and uses 21 GiB of disk memory.
Data sets used in experiment 2.
MNIST [28]. This dataset consists of grayscale pictures of handwritten digits of size $28\times 28$. We first flatten the images into vectors of size $784$, and then we project those vectors into Hamming space by setting each entry to 00 if it corresponded to a white pixel, i.e. if the grayscale color was 00, and to $1$ otherwise. Note that this transformation preserves the dataset fairly well, as the background pixels in each image are perfectly white.
Anonymous Microsoft Web Data (MSWeb) [12]. Entries in this dataset are anonymous web user IDs and paths on the Microsoft website that they have visited. There are in total $294$ different paths. We represent each user by a point in $294$ dimensional Hamming space, where a point entry is equal to $1$ if they have visited the corresponding path.
Mushroom [34, 38]. This dataset consists of descriptions of samples of mushrooms of $23$ different species. Each mushroom has $22$ categorical features describing it. We one-hot encode each feature to map the descriptions into Hamming space, with the final number of dimensions being $116$.
@@ Appendix C Additional experiments and details
In this experiment the goal was to measure the impact of $\lambda$ on the probability of success of the algorithm. The setup is the same as in Experiment 2. All runs are done on the Random dataset. The results are on Figure 3.
@@
Finally, we investigated the efficiency gain that our adaptive algorithm achieved over random point sampling. The experimental setup is the same as in Experiment 2. Here we use the Random dataset. For both approaches, we continued rerunning them until a false negative query was found. The results are presented in Figure 5.
@@
Experiment 6: Comparison of defense mechanisms.
In this experiment we compare two different defence mechanisms based on ideas from differential privacy [10] and distance estimation [14].
Theorem 1.3. of [14] presents a general robustification mechanism which, when applied to then LSH of [22], says that it can be made robust with blow-ups of $O(d\log 1/\delta)$ for memory and $O(\log 1/\delta)$ for query time, where $\delta$ is the failure probability per adaptive query. This is achieved by creating $O(d\log 1/\delta)$ copies of plain LSH. Querying is done by selecting $O(\log 1/\delta)$ of those LSH data structures, relaying the query to them and then choosing any of the returned points as the answer, or $\bot$ if all of the data structure failed to find a point.
As a side note, the original ADE construction of [14] would also allow to solve the ANN problem. We don’t study it in this experiment for two reasons: it would require time per query $\Omega(n)$, which is too slow in our parameter setting, and it is based upon linear sketches, against which our adaptive adversary is not effective.
The generic differentially private conversion of [10] is similar. For a generic estimation data structure, to be able to process $T$ queries against an adaptive adversary, they create $\widetilde{O}(\sqrt{T})$ copies of the data structure. For each query, they select $\widetilde{O}(1)$ copies of the data structure and aggregate their output using a differentially private median. Although their construction is not generally applicable to ANN, as it is a search problem, it would nevertheless still be effective against our adversary, as for the latter it is not important which exact point was returned, but only whether anything was returned at all.
[…]
The results are depicted on Figures 6 and 7. On both figures, plots annotated with both $\lambda$ and “q.s.” correspond to one of the two robustified variation accordingly. $\lambda$ represents the number of copies used, and “q.s.” stands for “query samples”, and represents the number of data structures randomly sampled to answer each query. Each used copy is build with the same parameters as in experiment $2$, and $\lambda=1$. The plots annotated with only $\lambda$ are the plain LSH constructions with different values of $\lambda$. The idea is that the plots with the same value of $\lambda$ have the same number of initialized hash functions. The used dataset is Random.
Both figures contain 4 subplots. The first subplot represents the probability that our adaptive adversary reports that it has found a false negative query at a desired target distance. Note that here, unlike in our previous experiments, the output of each query is non-deterministic even after the randomness of the hash functions is fixed. Hence we also test how “good” of a false negative the reported point is. For each found point, we repeatedly query it to the data structure $100$ times, and depict on the other three subplots the probability of the points discovered by the adversary to still be a negative query for at least $90\%$, $50\%$ and $10\%$ of the repeated queries.
As you can see from the Figure 6, the defense based on [14] is performing better than querying the plain LSH with the same and even often greater amount of hash functions. Similarly, the probability that the discovered false negative remain false negatives in at least $50\%$ or $90\%$ of cases is fairly low.
On the other hand, the results for the algorithm based on differential privacy on Figure 7 show that the probability of reporting a false negative is considerably higher than for a plain LSH even for low values of target distance. This is due to the aggregation mechanism that we use, as it has a chance of reporting nothing even if all sampled data structures returned a positive answer. However, this also means that the adversary has less chances to learn the underlying data structure, which we can see in the fact that the probability of that a repeated query would be a negative one in at least $10\%$ of cases being below $0.5$.
Based on this, it appears that unless in your application you are prepared to tolerate a significantly higher chance of failure for each query, it is better to use the construction based on [14]. Otherwise, a differentially private construction would leak less information to the adversary. Note that both constructions are still computationally faster than directly querying each data structure.
@@
Figure 6: Performance of robustification based on [14]. First subplot depicts probability that our adaptive adversary finds a false negative query versus the desired distance to that query from origin. The rest of subplots depict probabilities that, if such a query was found, if it was queried again it would still be a false negative query with probability at least $90\%$, $50\%$ or $10\%$. In the legend, plots annotated with both $\lambda$ and “q.s” correspond to the robust algorithm, with “q.s.” standing for “query samples”, the number of data structures sampled to answer each query, and $\lambda$ representing the total number of used data structures. The rest are the plain LSH construction with different values of $\lambda$.
Figure 7: Performance of robustification based on differential privacy [10]. First subplot depicts probability that our adaptive adversary finds a false negative query versus the desired distance to that query from origin. The rest of subplots depict probabilities that, if such a query was found, if it was queried again it would still be a false negative query with probability at least $90\%$, $50\%$ or $10\%$. In the legend, plots annotated with both $\lambda$ and “q.s” correspond to the robust algorithm, with “q.s.” standing for “query samples”, the number of data structures sampled to answer each query, and $\lambda$ representing the total number of used data structures. The rest are the plain LSH construction with different values of $\lambda$.